OpenStudy (anonymous):

solve the system of equations -3x+3y-15z=-12, x-y+5z=4, 2x-2y+10z=8

OpenStudy (anonymous):

1) -3x+3y-15z=-12 2) x-y+5z=4 3) 2x-2y+10z=8 2 * 1) 6 * 2) 3 * 3) -6x + 6y - 30z = -24 6x - 6y + 30z = 24 6x - 6y + 30z = 24 all the same line. so I would think that the answer is just one of the lines